right wrong
Pick an answer to see how members did
Log this miss in the free GMAT error log
Get the error logGet one like this every day: the free 60-day plan
Start the planA. 0.064
B. 0.36
C. 0.64
D. 0.784
E. 0.936
OA E
Source: Magoosh
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
right wrong
Pick an answer to see how members did
Log this miss in the free GMAT error log
Get the error logGet one like this every day: the free 60-day plan
Start the planGiven that the probability is 0.6 that an “unfair” coin will turn up tails on any given toss, the probability that it turns head = 1 – 0.6 = 0.4.BTGmoderatorDC wrote: ↑Sun Apr 05, 2020 5:38 pmThe probability is 0.6 that an “unfair” coin will turn up tails on any given toss. If the coin is tossed 3 times, what is the probability that at least one of the tosses will turn up tails?
A. 0.064
B. 0.36
C. 0.64
D. 0.784
E. 0.936
OA E
Source: Magoosh
ASIDE---------------BTGmoderatorDC wrote: ↑Sun Apr 05, 2020 5:38 pmThe probability is 0.6 that an “unfair” coin will turn up tails on any given toss. If the coin is tossed 3 times, what is the probability that at least one of the tosses will turn up tails?
A. 0.064
B. 0.36
C. 0.64
D. 0.784
E. 0.936
OA E
Source: Magoosh
Since the probability of obtaining tails on one toss is 0.6, then the probability of obtaining heads on one toss is 1 - 0.6 = 0.4.BTGmoderatorDC wrote: ↑Sun Apr 05, 2020 5:38 pmThe probability is 0.6 that an “unfair” coin will turn up tails on any given toss. If the coin is tossed 3 times, what is the probability that at least one of the tosses will turn up tails?
A. 0.064
B. 0.36
C. 0.64
D. 0.784
E. 0.936
OA E
Source: Magoosh
